OpenStudy (anonymous):

A website that sells songs for downloading increased its price per song from $0.99 to $1.29. Macy spent $15.36 downloading songs during the month of the price increase. She downloaded 4 more songs at $0.99 than at $1.29. What is the exact number of songs Macy downloaded at the $0.99 price?

Mehek (mehek14):

x = y + 4 0.99x + 1.29y = 15.36

Mehek (mehek14):

0.99x + 1.29y = 15.36 0.99(y+4)+ 1.29y= 15.36

Mehek (mehek14):

0.99y + 3.96 + 1.29y = 15.36 2.28y + 3.96= 15.36

Mehek (mehek14):

solve for y 2.28y + 3.96= 15.36 2.28y = 11.40 y = 5

Mehek (mehek14):

substitute y x = y + 4 x = 9

Mehek (mehek14):

@onexraygirl you there?

OpenStudy (anonymous):

so sorry.. my computer started acting weird so I rebooted. Thank you for your help! It is for my daughter...I am helping her with her review for a final and I am NOT good with word probs. If I can get it set up then I can figure it out....but not good with setting it up. I love that you gave step by step instructions...that makes it easier for me to explain to her. Thanks again!

Mehek (mehek14):

your very welcome :)